OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form. Answer y = -2x + 14 y = -2x + 5 y = -2x + 9 y = -2x + 6

OpenStudy (anonymous):

slope intercept form is displayed as y=mx+b m=slope, and b=your y-intercept

OpenStudy (anonymous):

rearrange your equation to fit this form. to do this you need to distribute your -2 on the right hand side, then add 4 to both sides to isolate your y variable.

OpenStudy (anonymous):

So... C?

OpenStudy (anonymous):

y-4=-2(x-5) with your -2 distributed it would look like y-4=-2x+10 then when you add 4 to each side to isolate your y variable you get y=-2x+14

OpenStudy (anonymous):

how did you get C?

OpenStudy (anonymous):

HAha. IDK!